OpenStudy (anonymous):

Find the equation of a parabola with focus (1,-2), and vertex (-3,-2)

OpenStudy (anonymous):

Not sure. :c

OpenStudy (anonymous):

I need to find the equation

OpenStudy (anonymous):

http://www.mathwords.com/f/focus_parabola.htm

OpenStudy (anonymous):

(y+2)^2=4*4(x+3)

OpenStudy (anonymous):

How'd you do that o:

OpenStudy (anonymous):

I'm a math noob, sorry. o;

OpenStudy (anonymous):

It's okay. xD

OpenStudy (anonymous):

equation of a parabola is y^2 = 4 a x......here a is the distance b/n focus and vertex nd the above eqn is wen verex is @ origion...so u got to shift the vertex to the given point

OpenStudy (anonymous):

*origin

OpenStudy (anonymous):

O: Ohoh. Okay. Tysm. <33